Hi! I was wondering if anyone could walk me through how to do this problem?
 Write a quadratic function given the following:
 (-1,6), (1,4), (2,9)
 y=
 Thanks!
 Answer by ankor@dixie-net.com(22740)
       
 Write a quadratic function given the following: (-1,6), (1,4), (2,9)
 :
 Using the form; ax^2 + bx + c = y; we will write three equations:
 eq1: x=-1; y=6, substitute for x and y
 -1^2a + -1b + c = 6
 1a - 1b + c = 6
 or just
 a - b + c = 6
 :
 eq2: x=1; y=4
 1^2a + 1b + c = 4
 a + b + c = 4
 :
 eq3: x=2, y=9
 2^2a + 2b + c = 9
 4a + 2b + c = 9
 :
 We will solve these equations by elimination
 Subtract eq 1 from eq 2, eliminate a and c
 a + b + c = 4
 a - b + c = 6
 -----------------subtraction eliminate a and c
 0 +2b + 0 = -2
 b = -1
 :
 Substitute -1 for b in the 1st equation
 a - (-1) + c = 6
 a + 1 = c = 6
 a + c = 6 - 1
 a + c = 5
 :
 do the same in the 3rd equation
 4a + 2(-1) + c = 9
 4a - 2 + c = 9
 4a + c = 9 + 2
 4a + c = 11
 :
 Using these two equations:
 4a + c = 11
 a  + c = 5
 -------------Subtraction eliminates c, find a
 3a = 6
 a = 2
 :
 Find c using a + c = 5
 2 + c = 5
 c = 5 - 2
 c = 3
 :
 Our equation: y = 2x^2 - x + 3
 :
 You can check our solution, substitute for x and ensure that it equals y for each given coordinate
 :
 Did you understand what we did here? Any questions?
 
 | 
  
 | 
 |